# Towards Efficient PCSEL Design: A Data-Driven Approach for Design Insights

### v1

@@
Towards Efficient PCSEL Design: A Data-Driven Approach for Design Insights
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
We present a data-driven design approach for photonic crystals to achieve high efficiency in photonic crystal surface-emitting lasers (PCSELs). By discretizing the photonic crystal structure into a grid, we enable the generation of arbitrary lattice designs. Multiple fully connected layers combined with a position embedding module extract essential features from the photonic crystal designs, while coupled-wave theory (CWT) is used to evaluate the efficiency based on the ratio of surface-emitting to edge-emitting resonant modes. The model achieves high prediction accuracy, with Pearson correlation coefficients of 0.88 for SEE and 0.78 for the log-transformed Q. Additionally, we perform Shapley value analysis to identify the most important Fourier coefficients, providing insights into the factors that impact the performance of PCSEL designs. Our work speeds up the design process and offers valuable guidance for optimizing high-performance PCSELs, supporting the development of fully photonic design automation (PDA).
1 Introduction
Photonic crystal surface-emitting lasers (PCSELs) have attracted significant interest due to their potential across a wide range of applications, including light detection and ranging (LIDAR) systems, laser processing, and optical communications [1, 2, 3, 4]. These devices offer the unique advantage of operating in a surface-emitting configuration, which provides several benefits over traditional semiconductor lasers, such as edge-emitting lasers (EELs) and vertical cavity surface-emitting lasers (VCSELs). These benefits include high output power, excellent beam quality, an ultra-small divergence angle, and single-mode operation [1, 5, 6]. Despite these advantages, traditional methods for designing photonic crystals often rely on trial-and-error or parameter sweep approaches [7, 8, 9, 4]. These methods can be both time-consuming and computationally expensive, making it difficult to efficiently explore the vast design space of PCSELs.
Data-driven approaches have gained traction in the field of photonics, as they provide a more efficient alternative for design optimization [10, 11]. These methods leverage machine learning models to explore complex design spaces and identify optimal solutions without the need for exhaustive simulations. By applying these techniques, the design process can be significantly accelerated, and the computational burden reduced [12, 13]. However, most existing approaches focus on general optimization rather than providing insights into the specific features of photonic crystals that drive performance, which limits their applicability to high-efficiency PCSEL design.
@@ 1 Introduction
2 Methods
2.1 Photonic crystal simulation
@@
For detailed calculations of the elements in the matrices of Eq. (2), Eq. (3) and Eq. (4), please refer to Ref. [16, 18].
@@
2.2 Neural network architecture
@@
The network consists of four fully connected layers with 1024, 512, 256, and 64 neurons, respectively. Each layer is followed by a Parametric ReLU (PReLU) activation function, which introduces nonlinearity and aids in learning complex patterns. Finally, an output layer with a single neuron predicts $SEE$ and $\log{Q}$. Note that a sigmoid activation function is applied to the output for $SEE$ to ensure its value is constrained within the range $(0,1)$.
@@
2.3 Generation of train and test dataset
@@
The simulation function is subsequently called with the generated dielectric array and a solver instance. This function performs optical simulations based on coupled-wave theory (CWT) and finite element method (FEM) analysis, and returns key performance metrics such as the $SEE$ and quality factor $Q$, and other relevant parameters. We adopt the epitaxial structure outlined in Table 1, with the Bragg wavelength fixed at 980 nm. To ensure consistency with the finite-size PCSELs, the photonic crystal’s lateral dimension $L$ is chosen to be as close as possible to 200 $\mu m$, while maintaining an integer multiple of the crystal period. The photonic crystal unit cells are discretized using a $17\times 17$ grid to capture the structural details accurately. The results are then written to a NPZ file [21], which serves as the dataset for further model training and testing. This iterative process continues until an external stop signal is detected, thereby building a comprehensive dataset of photonic crystal designs. Finally, our dataset consists of 200k simulation results, ensuring a diverse and representative set of samples for training robust machine learning models.
@@
2.4 Shaply values analysis
3 Results and discussion
3.1 Model training and testing
@@
As shown in Table 2 these results demonstrate that the neural network model effectively captures the key performance characteristics of PCSEL designs. The Pearson correlation coefficients which means the linear relationship between predicted and actual values, are 0.780 for $SEE$ and 0.887 for $\log{Q}$. These high correlation values indicate that the model has strong predictive capability. The relatively low RMSE (Root of Mean Square Error) values further suggest that the model produces accurate estimates with a minimal deviation from the simulated values. We also calculated the coefficient of determination ($R^{2}$) to evaluate the proportion of variance in the data that is explained by the model. An $R^{2}$ value closer to 1 indicates that the model explains a larger portion of the variance in the data, confirming its predictive power. In our results, the $R^{2}$ values of 0.595 for $SEE$ and 0.784 for $\log{Q}$ confirm that the model explains a substantial portion of the variance in the data. These findings highlight the potential of data-driven approaches for accelerating the optimization of PCSEL designs, reducing reliance on computationally expensive simulations while maintaining high prediction accuracy.
3.2 Shaply values analysis
@@
To gain insights into the relative importance of different input features in determining the efficiency of PCSEL designs, we conducted a Shapley value analysis (on 500 randomly selected samples from the test dataset). The results of this analysis are shown in Figure 3, which illustrates the contribution of each Fourier coefficient to the predicted $SEE$ and $\log{Q}$. The Shapley values provide a quantitative measure of the impact of each input parameter on the model’s output, revealing the most influential features in the photonic crystal structure.
@@
$\xi^{R}_{2,0}$ and $\xi^{R}_{0,2}$ also exhibit a significant impact on both $SEE$ and $\log{Q}$. It can be observed that when $\xi^{R}_{2,0}$ and $\xi^{R}_{0,2}$ are relatively low, both SEE and $\log{Q}$ tend to decrease.
Other Fourier coefficients also contribute to the results, but their influence is less straightforward, indicating that there are complex interactions between them.
Table 3: Summary of the influence of Fourier coefficients on surface-emitting efficiency ($SEE$) and the quality factor ($\log{Q}$). The symbols "+" and "-" indicate positive and negative correlations, respectively. The number of "+" or "-" symbols represents the magnitude of the influence, with more symbols indicating a stronger effect on $SEE$ and $\log{Q}$. "00" indicates no significant impact.
@@
$\xi^{R}_{2,0}$ and $\xi^{R}_{0,2}$ $+$ $+$
$\xi^{R}_{1,1}$ $+$ 00
$\xi^{R}_{1,-1}$- --
4 Conclusion
@@
[…]
Funding This research was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 62174144, the Shenzhen Science and Technology Program under Grants No. JCYJ
20220818102214030, No. KJZD20230923115114027, No. JSGG20220831093602004, No. KJZD
20240903095602004, the Guangdong Key Laboratory of Optoelectronic Materials and Chips under Grant No. 2022KSYS014, the Shenzhen Key Laboratory Project under Grant No. ZDSYS201603311644527, the Longgang Key Laboratory Project under Grants No. ZSYS2017003 and No. LGKCZSYS2018000015, and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300701), Hefei National Laboratory, Hefei 230088, China.
[…]
Acknowledgment The authors would thank Ms. Floria Chen and Ms. Chuanyi Yang for their mentorship and revision tips. The authors would also thank Mr. Qi Xin and Mr. Sixuan Mao for his suggestions on frameworks of artificial neural networks.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data availability Parts of Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request. The dataset for the NN model is available Ref. [21].
[…]
Supplemental document A supplemental document must be called out in the back matter so that a link can be included. For example, “See Supplement 1 for supporting content.” Note that the Supplemental Document must also have a callout in the body of the paper.

### v3

@@
Abstract
We present a data-driven design approach for photonic crystals to achieve high efficiency in photonic crystal surface-emitting lasers (PCSELs). By discretizing the photonic crystal structure into a grid, we enable the generation of arbitrary lattice designs. Multiple fully connected layers combined with a position embedding module extract essential features from the photonic crystal designs, while coupled-wave theory (CWT) is used to evaluate the efficiency (based on the ratio of surface-emitting to edge-emitting resonant) and quality factor Q. We introduce the Neural Networks (NNs) model to evaluate the structures, and to find a better performance design according to the evaluation result. The model achieves high prediction accuracy, with Pearson correlation coefficients of 0.780 for SEE and 0.887 for the log-transformed Q. Additionally, we perform Shapley value analysis to identify the most important Fourier coefficients, providing insights into the factors that impact the performance of PCSEL designs. Our work speeds up the design process and offers valuable guidance for optimizing high-performance PCSELs, supporting the development of fully photonic design automation (PDA).
Introduction
Photonic crystals are structures with periodic dielectric materials that create photonic band structures, enabling the control of light propagation precisely. These structures have been widely used in various applications, including waveguides, filters, and lasers [1, 2, 3]. Photonic crystal surface-emitting laser (PCSEL), an important application of photonic crystals, has attracted significant interest due to their potential across a wide range of applications, including light detection and ranging (LIDAR) systems, laser processing, and optical communications [4, 5]. These devices offer the unique advantage of operating in a surface-emitting configuration, which provides several benefits over traditional semiconductor lasers, such as edge-emitting lasers (EELs) and vertical cavity surface-emitting lasers (VCSELs). These benefits include high output power, excellent beam quality, an ultra-small divergence angle, and single-mode operation [6, 7, 8]. Despite these advantages, traditional methods for designing photonic crystals often rely on trial-and-error or parameter sweep approaches [9, 10, 11, 12, 13]. These methods can be both time-consuming and computationally expensive, making it difficult to efficiently explore the vast design space of PCSELs.
Data-driven approaches have gained traction in the field of photonics, as they provide a more efficient alternative for design optimization [14, 15]. These methods leverage machine learning models to explore complex design spaces and identify optimal solutions without the need for exhaustive simulations. By applying these techniques, the design process can be significantly accelerated, and the computational burden reduced [16, 17]. However, most existing approaches focus on general optimization rather than providing insights into the specific features of photonic crystals that impact PCSEL, which limits their applicability to high-efficiency PCSEL design.
@@ Introduction
Methods
Photonic crystal simulation
@@ Methods
For detailed calculations of the Eq. (2), Eq. (3) and Eq. (4), please refer to Ref. [20, 22].
@@
Although the partial differential equation (PDE) analysis based on CWT is computationally fast (requiring only a few seconds), it becomes a significant bottleneck when performing large-scale parameter optimization and analysis. Consequently, to address this challenge and speed up the evaluation process, we propose the use of neural networks to predict $SEE$ and $\log{Q}$, bypassing the need for extensive numerical computations. By training a neural network model on the simulated data, we can achieve rapid predictions of $SEE$ and $\log{Q}$, significantly reducing the time required for optimization.
Neural network architecture
@@
The network consists of four fully connected layers with 1024, 512, 256, and 64 neurons, respectively. Given that the matrix is complex, it is logical to concatenate real and Imaginary latent matrix, using a Multi-Layer Perceptron (MLP) to learn the interaction. In this context, NNs can be defined as
$z_{1}$ $=W_{0}^{T}x+b_{0},$ (9)
$\phi(z_{1})$ $=a(W_{1}^{T}z_{1}+b_{1}),$ (10)
$\dots$,… , (11)
$\phi(z_{4})$ $=a(W_{4}^{T}z_{4}+b_{4}),$ (12)
$y$ $=\sigma(\phi(z_{4})),$ (13)
where $W_{x}$ and $b_{x}$ denoted as layer weights, $a$ is Parametric ReLU (PReLU) activations and $\sigma$ is a sigmoid function in training SEE to ensure its value is constrained within the range $(0,1)$ and $\sigma$ is identity transformation in training $\log Q$.
@@
Generation of dataset
@@
The simulation function is subsequently called with the generated dielectric array and a solver instance. This function performs optical simulations based on coupled-wave theory (CWT) and finite element method (FEM) analysis, and returns key performance metrics such as the $SEE$ and quality factor $Q$, and other relevant parameters. We adopt the epitaxial structure outlined in Table 1, with the Bragg wavelength fixed at 980 nm. To ensure consistency with the finite-size PCSELs, the photonic crystal’s lateral dimension $L$ is chosen to be as close as possible to 200 $\mu m$, while maintaining an integer multiple of the crystal period. The photonic crystal unit cells are discretized using a $17\times 17$ grid to capture the structural details accurately. The results are then written to a NPZ file (see supplementary information), which serves as the dataset for further model training and testing. This iterative process continues until an external stop signal is detected, thereby building a comprehensive dataset of photonic crystal designs. Finally, our dataset consists of 200k simulation results, ensuring a diverse and representative set of samples for training robust machine learning models.
@@
Results and discussions
Model training and testing
@@
As shown in Table 2 these results demonstrate that the neural network model effectively captures the key performance characteristics of PCSEL designs. The Pearson correlation coefficients which means the linear relationship between predicted and actual values, are 0.780 for $SEE$ and 0.887 for $\log{Q}$. These high correlation values indicate that the model has strong predictive capability. The relatively low RMSE (Root of Mean Square Error) values further suggest that the model produces accurate estimates with a minimal deviation from the simulated values. We also calculated the coefficient of determination ($R^{2}$) to evaluate the proportion of variance in the data that is explained by the model. A $R^{2}$ value closer to 1 indicates that the model explains a larger portion of the variance in the data, confirming its predictive power. In our results, the $R^{2}$ values of 0.595 for $SEE$ and 0.784 for $\log{Q}$ confirm that the model explains a substantial portion of the variance in the data. These findings highlight the potential of data-driven approaches for accelerating the optimization of PCSEL designs, reducing reliance on computationally expensive simulations while maintaining high prediction accuracy.
Global optimization of photonic crystal design
After establishing the predictive capabilities of our trained neural network models for surface-emitting efficiency ($SEE$) and quality factor ($Q$), we leverage these models to search for a photonic crystal design that minimizes $\left|SEE_{Target}-SEE\right|$ while ensuring $Q>10^{4}$ to maintain a low operational threshold.
To achieve this, we define a loss function:
$\mathcal{L}=\left|SEE_{Target}-SEE\right|+\lambda\max(0,10^{4}-Q),$ (14)
where $SEE_{Target}$ is the desired surface-emitting efficiency, and $\lambda$ (is setted to 5e-5) is a penalty coefficient that enforces the quality factor constraint. A larger penalty is applied when $Q$ falls below the required threshold, ensuring that solutions with desired $SEE$ and sufficiently large $Q$ are favored.
We employ a global optimization algorithm, specifically the DIRECT [25, 26] (DIviding RECTangles) method, to efficiently explore the parameter space. The optimization variables include the positions, sizes, and orientations of the holes in the photonic crystal unit cell, as well as the filling factor. By utilizing the trained models for $SEE$ and $Q$, the optimization process is significantly accelerated, as each evaluation bypasses the need for relatively computationally expensive coupled-wave theory (CWT) simulations. To validate the optimal design obtained via the model-driven optimization, we perform full physical simulations using the CWT and FEM solvers. A comparison between predicted and simulated results further confirms the reliability of the neural network models in guiding the optimization process. The optimized structure is summarized in Table 3. The final validation results demonstrate strong agreement with our model predictions. The entire optimization process involved over 16,000 function evaluations, with a total runtime of approximately only 2.2 minutes. This significantly reduced the computational burden compared to direct CWT simulations. This demonstrates the effectiveness of combining machine learning-based surrogate models with global optimization algorithms for rapid and accurate photonic crystal design.
Table 3: Optimized photonic crystal structure for surface-emitting efficiency ($SEE$) and quality factor ($Q$). The dielectric constant distributions correspond to different $SEE_{Target}$ values, and both predicted and validated performance metrics are listed.
$SEE_{Target}$ $1.000$ $0.500$ $0.000$
Dielectric constant distribution [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Results $SEE$ $Q$ $SEE$ $Q$ $SEE$ $Q$
Predicted 0.889 10001.9 0.500 18643.2 0.034 13508.0
Validated 0.867 9524.2 0.477 19015.2 0.017 16666.2
Number of calls 16015 16135 16007
Total runtime 130.4s 143.1s 131.2s
This approach demonstrates the potential of machine learning models not only in predicting photonic crystal performance but also in guiding global optimization for high-efficiency laser design.
Shaply values analysis
@@
To gain the relative impacts of different input features in determining the performance of PCSEL designs, we conducted a Shapley value analysis (on 500 randomly selected samples from the test dataset). The results of this analysis are shown in Figure 3, which illustrates the contribution of each Fourier coefficient to the predicted $SEE$ and $\log{Q}$. The Shapley values provide a quantitative measure of the impact of each input parameter on the model’s output, revealing the most influential features in the photonic crystal structure.
@@
We summarized the impact of part of different Fourier coefficients on $SEE$ and $\log{Q}$ in Table 4. Other Fourier coefficients also contribute to the results, but their influence is less straightforward, indicating that there are complex interactions between them.
Table 4: Summary of the influence of Fourier coefficients on surface-emitting efficiency ($SEE$) and the quality factor ($\log{Q}$). The symbols "+" and "-" indicate positive and negative correlations, respectively. The number of "+" or "-" symbols represents the magnitude of the influence, with more symbols indicating a stronger effect on $SEE$ and $\log{Q}$. "\" indicates that the feature does not has direct impact on the corresponding result.
@@
$\xi^{R}_{1,1}$ $+$ \
$\left|\xi^{R}_{1,1}-\overline{\xi^{R}_{1,1}}\right|$ and $\left|\xi^{R}_{-1,1}-\overline{\xi^{R}_{-1,1}}\right|$-
$\xi^{R}_{2,0}$ and $\xi^{R}_{0,2}$ $+$
$\xi^{R}_{4,0}$- \
Conclusion
@@
Acknowledgements
This research was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 62174144, the Shenzhen Science and Technology Program under Grants No. JCYJ20220818102214030, No. KJZD20230923115114027, No. JSGG20220831093602004, No. KJZD20240903095602004, the Guangdong Key Laboratory of Optoelectronic Materials and Chips under Grant No. 2022KSYS014, the Shenzhen Key Laboratory Project under Grant No. ZDSYS201603311644527, the Longgang Key Laboratory Project under Grants No. ZSYS2017003 and No. LGKCZSYS2018000015, and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300701), Hefei National Laboratory, Hefei 230088, China.
The authors would thank Ms. Floria Chen and Ms. Chuanyi Yang for their mentorship and revision tips. The authors would also thank Mr. Qi Xin and Mr. Sixuan Mao for their suggestions on frameworks of artificial neural networks.
Author contributions statement
H.H. conceived the experiment and conducted the majority of the manuscript writing. H.H. and Z.X. conducted the experiment and analyzed the results. Z.X. wrote the section on the principles of the neural network and prepared Figure 1. Z.Z. provided guidance and funding support. All authors reviewed the manuscript.
Additional information
Competing interests The authors declare no conflicts of interest.